Litho status report 1/5

Gary Sosa gsosa at stanford.edu
Wed Jan 5 08:07:27 PST 2011


Good morning All.....

     Here is the Litho morning status report.

ASML Stepper- Down for annual PM and reticle table problem. FSE to 
continue work today

EVG Spraycoater- Down. Computer OK. Received and will install 
registration code for application software
                                  and continue to bring up tool to 
operational state

Laurell Spinner- Down. Spinner will be delivered today, late AM. 
Installation/testing to follow

SVG Coat 1- Down. Recipe Manager computer being repaired by IT Department

Blue M Ovens 110deg and 150deg- Temperature controllers replaced with 
new digital controllers over
                                                                 the 
shutdown to fix the temperature drift problems

Karl Suss Aligner 1- PM scheduled for today

Karl Suss Aligner 2- Verification of BSA microscope to complete PM. Post 
PM testing/adjusting to improve resolution

SVG Coaters- Note: SPR 3617 photoresist moved from Track 1(front) to 
Track 2(back). Associated recipes
                                     created in Recipe Manager and tested.

                                     SPR 220-3 photoresist moved from 
Track 2(back) to Track 1(front). Basic recipe created in Recipe
                                                          Manager, with 
all needed variations to be created after computer repairs are completed.

    Please check back in Coral for latest tool updates and current 
equipment status.

      Thanks...  The Litho Team



*/Photolithography Area Status/*

	

	

*/Wednesday, January 05, 2011/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

*/Exposure Tools/**//*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*DOWN*

	

Annual PM in Progress

	

1

EVAlign

	

*UP*

	

	

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

Scheduled PM today

	

Karl Suss 2

	

*PROBLEM*

	

Post PM testing

	

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*DOWN*

	

Will not boot up

	

4

Headway

	

*UP*

	

	

Laurel

	

*DOWN*

	

Tool out for PM

	

3

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Down*

	

Computer problem

	

2

3617 Process

	

*Down*

	

	

220-7 Process

	

*Down*

	

Optimizing process

	

SVGCoat2

	

3612 Process

	

*Up*

	

Intermittant transfer problem

	

955-0.7 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*UP*

	

Controller replaced during shutdown

	

Blue M 150 C

	

*UP*

	

Controller replaced during shutdown

	

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

-11 stage error

	

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*PROBLEM*

	

Left flag did not pull

	

KSBonder

	

*PROBLEM*

	

Leaking nitrogen hose

	

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

	

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20110105/8dd19e41/attachment.html>


More information about the litho mailing list